# A note on maximal plane subgraphs of the complete twisted graph containing perfect matchings

### v1

@@ 3 Main result
By definition, for $i=1,2,\ldots,k$, $M_{i}$ is obtained from $M_{i-1}$ by exchanging the edges of an alternating non-crossing cycle with four edges. Therefore, for $i=1,2,\ldots,k$, $M_{i-1}\cup M_{i}$ is a set of pairwise non-crossing edges of $T_{2m}$ that can be extended to a maximal plane subgraph $S_{i}$ of $T_{2m}$.
Let $S_{0}=S$ and $S_{k+1}=R$. Notice that for $i=0,1,\ldots,k$, $S_{i}$ and $S_{i+1}$ are maximal plane graphs of $T_{2m}$ that contain $M_{i}$. By Theorem 2, the graph $MP(T_{n},M_{i})$ contains a path joining $S_{i}$ and $S_{i+1}$. Clearly $MP(T_{n},M_{i})$ is a subgraph of $MP_{\cal{M}}(T_{2m})$ for each $i=0,1,\ldots,k$. Therefore the graph $MP_{\cal{M}}(T_{2m})$ contains a path joining $S$ and $R$. ∎

### v2

@@ 3 Main result
By definition, for $i=1,2,\ldots,k$, $M_{i}$ is obtained from $M_{i-1}$ by exchanging the edges of $M_{i}$ and $M_{i-1}$ of an alternating non-crossing cycle with four edges. Therefore, for $i=1,2,\ldots,k$, $M_{i-1}\cup M_{i}$ is a set of pairwise non-crossing edges of $T_{2m}$ that can be extended to a maximal plane subgraph $S_{i}$ of $T_{2m}$.
Let $S_{0}=S$ and $S_{k+1}=R$. Notice that for $i=0,1,\ldots,k$, $S_{i}$ and $S_{i+1}$ are maximal plane graphs of $T_{2m}$ that contain $M_{i}$. By Theorem 3, the graph $MP(T_{2m},M_{i})$ contains a path joining $S_{i}$ and $S_{i+1}$. Clearly $MP(T_{2m},M_{i})$ is a subgraph of $MP_{\cal{M}}(T_{2m})$ for each $i=0,1,\ldots,k$. Therefore the graph $MP_{\cal{M}}(T_{2m})$ contains a path joining $S$ and $R$. ∎